If P=5/6 and n=125 find P( x≤ 100)
using normal distribution on TI-83 exact steps

Answer by stanbon(75887)

If P=5/6 and n=125 find P( x≤ 100)
using normal distribution on TI-83 exact steps
----
u = np = 125(5/6) = 104.16
s = sqrt(npq) = sqrt(104.16*(1/6)) = 4.1667
------
P(x <= 100) = P(0<= x <=100.5)
----
z(0) = (0-104.16)/4.1667 = -24.9984
z(105) = (100.5-104.16)/4.1667 = -0.8784
----
P(0<= x <=100.5) = P(-25<= z <= -0.8784) = normalcdf(-25,-0.8784) = 0.1899
===================================
Cheers,
Stan H.
=========================